# Application of Non-Linear Noise Regression in the Virgo Detector

### v3

@@ 5 Summary and Conclusions
  • Data access: To address these high computational demands, the VirgoTool Python package [^6] https://anaconda.org/conda-forge/pythonvirgotools was identified as a potential solution for reducing data reading times. However, integration challenges within our cluster’s environment limited its utility. We suggest enhancing the gwpy library with VirgoTool’s data handling capabilities to improve computational efficiency in future studies.
6 Acknowledgements
@@
R.W.K were supported by the Axe “ondes gravitationnelles/multimessager” de l’Observatoire de la Côte d’Azur, Boulevard de l’Observatoire, F-06304 Nice. M.S. and M.W.C acknowledge the support from the National Science Foundation with grant numbers PHY-2308862 and PHY-2117997.
@@
The authors are very grateful to Alba Romero Rodríguez for her review on behalf of Virgo DRS and P&P, which has been useful in the improving of this paper. We would also like to thank Nathalie Besson for her valuable feedback, which enhanced the clarity and precision of the manuscript.
@@
The authors acknowledge the IGWN Computing Grid (CIT, LHO, LLO) for providing resources to realize our simulations. We acknowledge the Virgo DetChar and Virgo cluster at Cascina for providing resources that contributed to the research results reported within this paper. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility operated under contract No. DE-AC02-05CH11231 under project “Toward a complete catalog of variable sources to support efficient searches for compact binary mergers and their products.” This work used the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS). This work used the Extreme Science and Engineering Discovery Environment (XSEDE) COMET at SDSU through allocation AST200016 and AST200029. This material is based upon work supported by NSF’s LIGO Laboratory, which is a major facility fully funded by the National Science Foundation

### v4

@@ V Summary and Conclusions
  • Data access: To mitigate data read latency, we explored the VirgoTool Python package [^8] https://anaconda.org/conda-forge/pythonvirgotools. However, compatibility issues with our cluster limited its use. We recommend integrating VirgoTool’s data handling capabilities into the more widely adopted gwpy library to streamline data access and improve computational efficiency in future studies.
  • Configuration: Moving from LIGO to Virgo required only configuration changes, mainly in the choice of witness channels, frequency-band setup, and training parameters, while the DeepClean algorithm itself remained unchanged. This demonstrates its transferability.
VI Acknowledgements
@@
R.W.K were supported by the Axe “ondes gravitationnelles/multimessager” de l’Observatoire de la Côte d’Azur, Boulevard de l’Observatoire, F-06304 Nice.
R.W.K acknowledge ANR (French National Research Agency) for its support of the project ”Multi-messenger observations of the Transient Sky (MOTS)” no. ANR-22-CE31-0012.
M.S. and M.W.C acknowledge the support from the National Science Foundation with grant numbers PHY-2308862 and PHY-2117997. M.S. also acknowledges the support from Weinberg Institute for Theoretical Physics at the University of Texas at Austin.
@@
The authors thank Franco Carbognani for making the O3b Virgo data available on the CIT cluster for our work. We are also very grateful to Alba Romero Rodríguez for her review on behalf of Virgo DRS and LIGO P&P, which has been useful in the improving of this paper. We would also like to thank Nathalie Besson for her valuable feedback, which enhanced the clarity and precision of the manuscript.
@@
The authors acknowledge the International Gravitational-wave Network (IGWN) Computing Grid (CIT, LHO, LLO) for providing resources to realize our simulations. We acknowledge the Virgo DetChar and Virgo cluster at Cascina for providing resources that contributed to the research results reported within this paper. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility operated under contract No. DE-AC02-05CH11231 under project “Toward a complete catalog of variable sources to support efficient searches for compact binary mergers and their products.” This work used the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS). This work used the Extreme Science and Engineering Discovery Environment (XSEDE) COMET at SDSU through allocation AST200016 and AST200029. This material is based upon work supported by NSF’s LIGO Laboratory, which is a major facility fully funded by the National Science Foundation
APPENDIX
.1 Popular subtraction techniques
Wiener filtering—This is a fundamental filtering technique that is effective when the noise is stationary and noise coupling is linear. It aims at minimizing the mean square error (MSE) between estimated and true signals [21, 12]. Its robustness is evident in GW event analysis by the IGWN [22]. However, its performance becomes sub-optimal for non-stationary and non-linear noise driven by complex environmental and instrumental fluctuations. This highlights the need for adaptive noise reduction strategies in GW astronomy.
NonSENS—Non-Stationary Estimation of Noise Subtraction (NonSENS) employs an Infinite Impulse Response (IIR) filter to remove both stationary and non-stationary noise in GW detectors [14]. This method effectively addresses noise arising from slow interferometer movements, such as angular fluctuations, which pose challenges to traditional techniques. By utilizing witness signals to monitor noise and its modulation over time, NonSENS achieves stable, parametric subtraction of non-stationary noise. It leverages data science techniques, such as gradient-based optimization [31], to estimate the coupling coefficients, thereby overcoming the complexities that hinder standard noise reduction methods. Additionally, the analytical filter, with its limited number of coefficients, avoids the computational challenges and interpretability issues often associated with deep neural networks.
Adaptive Feed-Forward—During the O3 run, Advanced Virgo implemented an adaptive linear feed-forward technique to suppress 50 Hz noise originating from the electrical mains [20]. A phase signal from the Uninterruptible Power Supply (UPS) in the Central Building serves as a witness channel for this disturbance. The signal is filtered through a narrow-band resonant filter centered at 50 Hz and adjusted in amplitude (gain) and phase to match the coupling of the noise into the differential arm length (DARM) channel. The resulting correction is then injected into DARM, effectively cancelling the targeted line. This control is adaptive: the gain and phase are continuously updated in real time based on demodulation of DARM with respect to the UPS signal, allowing the system to track variations in the noise coupling path and maintain robust suppression of the 50 Hz line.